Your boss has asked you to evaluate the economic viability of refinancing a loan on your​ plant's process equipment. The original loan of​$700,000 was for 7 years. The payments are monthly and the nominal interest rate on the current loan is 9​%per year. As of the present​ time, your company has had the loan for 36 months. The new loan would be for the current balance​ (i.e. the balance at the end of the 36th month on the old​ loan) with monthly payments at a nominal interest rate of 3​%per year for 4 years. A​ one-time financing fee for the new loan is​ $20,000.Your​ company's MARR is 12​% per year on a nominal basis. Determine if the new loan is economically advantageous.

What is the present worth of the difference between the original financing plan and the new​ (proposed) financing plan?
Answer #1

old loan

loan amount = 700000

loan term = 7*12 = 84 months

loan int rate = 9% / 12 = 0.75% per month

monthly payment = 700000 * (A/P,0.75%,84) = 700000 * 0.016089 = 11262.35

Loan balance after 36th payment , 48 payments are left

Loan balance = 11262.35 * (P/A,0.75%,48) = 11262.35 * 40.184782 = 452575.27

New loan amount = 452575.27 + 20000 = 472575.27

New loan int rate = 3%/12 = 0.25% per month

New loan monthly payment = 472575.27 * (A/P,0.25%,48) = 472575.27 * 0.022134 = 10460.14

Savings per month = 11262.35 - 10460.14 = 802.21

As there is savings in loan payments, therefore new loan is beneficial

MARR = 12% = 12%/12 = 1% per month

PW of savings = 802.21 * (P/A,1%,48) = 802.21 * 37.973959 = 30463.26